OpenStudy (anonymous):

Which statement shows the commutative property of multiplication? A. 7 · (4 · x) = (7 · 4) · x B. 18 + 6y = 6y + 18 C. If 5h = 30, then 30 = 5h D. (d + 4) ·17 = 17 · (d + 4)

TheSmartOne (thesmartone):

Commutative Property of Multiplication \[a \times b =b \times a\]

OpenStudy (gorv):

a*b=b*a is commutaitve property

TheSmartOne (thesmartone):

Or \[a \times (b \times c)= (a \times b) \times c\]

TheSmartOne (thesmartone):

What do you think the answer is @Freedom101